OpenStudy (anonymous):

integrate ... x^3/sqrt(x^2-5)

myininaya (myininaya):

\[u=x^2-5=> du=2x dx \text{ and } u=x^2-5=>u+5=x^2\]

myininaya (myininaya):

\[\text{ write } x^3 \text{ as } x^2 \cdot x \text{ first}\]

OpenStudy (anonymous):

how do u do it? i think its other style...

myininaya (myininaya):

\[\int\limits_{}^{}\frac{x^2 \cdot x}{\sqrt{x^2-5}} dx\]

OpenStudy (anonymous):

yes i got that..

myininaya (myininaya):

\[\int\limits_{}^{}\frac{(u+5)}{\sqrt{u}} \frac{du}{2}\]

myininaya (myininaya):

\[\frac{1}{2} \int\limits_{}^{}u^\frac{-1}{2}(u+5) du\]

OpenStudy (anonymous):

why u+5/sqrt u??

myininaya (myininaya):

\[x^2=u+5; x dx =\frac{du}{2}; x^2-5=u\]

myininaya (myininaya):